Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of theRequired Required Required Required 2 3 Should the outside suppliers offer be accepted? Yes ONO < Required 1 Required 3 >Required Required Required Required 3 4 Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freGiven the new assumption in requirement 3, should the outside suppliers offer be accepted? Yes ONO
Answer #1

1) Differential analysis

Make Buy
Direct material 256000
Direct labor 192000
Variable manufacturing overhead 48000
Fixed manufacturing overhead (48000/3) 16000
Purchase cost (16000*35) 560000
Total relevant cost 512000 560000

Financial (disadvantage) = 512000-560000 = -48000

2) No

3) Differential analysis

Make Buy
Direct material 256000
Direct labor 192000
Variable manufacturing overhead 48000
Fixed manufacturing overhead (48000/3) 16000
Opportunity Cost 160000
Purchase cost (16000*35) 560000
Total relevant cost 672000 560000

Financial advantage = 672000-560000 = 112000

4) Yes